OpenStudy (anonymous):

Which of the following is a factor of 27x^3 – 64y3? 3x + 8y 9x^2 + 12xy + 16y^2 9x^2 + 24xy + 16y^2 None of the above

OpenStudy (anonymous):

Which of the following is a factor of 27x^3 – 64y^3? ****

OpenStudy (anonymous):

(3x-8y){(3x)^2 + (3x)(8y) + (8y)^2}

OpenStudy (anonymous):

\[x^3-y^3=(x-y)(x^2+xy+y^2) \]

OpenStudy (anonymous):

nathematicalpro.... none of the above right?

OpenStudy (anonymous):

because 8^2 = 64

OpenStudy (anonymous):

9x^2 + 12xy + 16y^2

OpenStudy (anonymous):

27x^3 - 64y^3 = (3x}^3 - (4y)^3 = [ 3x - 4y ][ (3x)^2 + (3x)(4y) + (4y)^2 ] // since a^3-b^3 = (a-b)(a^2+ab+b^2) = ( 3x - 4y )( 9x^2 + 12xy + 16y^2 ) therefore 9x^2 + 12xy + 16y^2 is a factor of 27x^3 – 64y3